# Exponential Convergence of Augmented Primal-dual Gradient Algorithms for Partially Strongly Convex Functions

### v1

@@ Abstract
We show that global exponential convergence for the augmented primal-dual gradient algorithms can be achieved for partially strongly convex functions. In particular, the objective function only needs to be strongly convex in the subspace satisfying the equality constraint and can be generally convex elsewhere provided the global Lipschitz condition for the gradient is satisfied. This condition also implies that states outside the equality subspace will be driven towards it by the augmented primal-dual algorithm exponentially fast. The analysis is then applied to distributed optimization, where the partially strong convexity can be relaxed to the restricted secant inequality condition, which is not necessarily convex. This work unifies global exponential convergence results for some existing centralized and distributed algorithms.
@@ I Introduction
The convergence of optimization algorithms has been an important research focus due to their wide range of applications. In particular, exponential convergence in continuous time, or linear convergence in discrete time, has attracted considerable attention recently. Researchers have proposed various conditions alternative to strong convexity of the objective function to ensure exponential/linear convergence, such as the Polyak-Łojasiewicz (PL) condition [1, 2], the error bound (EB) condition [3], the restricted secant inequality (RSI) condition [4], and the quadratic growth (QG) condition [5]. Notably, these conditions may not even require convexity. A detailed discussion of the relationships between these conditions can be found in [6] and references therein.
Optimization becomes more complex when constraints are introduced. The augmented Lagrange multiplier method for handling equality constraints was introduced by [7, 8], and has been extensively studied since then [9]. Exponential convergence of primal-dual gradient algorithms has also been widely studied in recent years [10, 11, 12, 13, 14]. Specifically, [10] and [11] address this issue for strongly convex functions with linear equality and inequality constraints, respectively. For general nonlinear inequality constraints, only the so-called semi-global exponential stability is obtained without considering constraints to be linearly independent [12]. Currently, it remains uncertain whether a global exponential stability condition can be established for (augmented) primal-dual gradient dynamics in problems involving general nonlinear inequality constraints. The work [14] extends exponential stability to non-smooth objective functions and general set constraints. Moreover, its application in distributed optimization is also extensively studied [15, 16, 17].
On the other hand, exponential convergence of primal-dual gradient algorithms by harnessing strong convexity from constraints is seldom considered. The most related work is [13], which addresses partially non-strongly convex problems but with decoupled objective functions and equality constraints. In this work, we aim to explicitly demonstrate how equality constraints can contribute to exponential convergence while relaxing the requirement for strong convexity. Our theoretical results are shown by utilizing the methods of integral quadratic constraints (IQCs) [18] and the Kalman-Yakubovich-Popov (KYP) lemma [19]. The proposed framework unifies and extends some existing results in both centralized and distributed optimization, providing a comprehensive understanding of exponential convergence in these contexts.
@@ II Preliminaries
$\min_{x\in\mathbb{R}^{n}}f(x)\quad\text{subject to }Tx=b$ (1)
@@
$f(x)$ is $l$-Lipschitz smooth, i.e., for some constant $l>0$, the following inequality holds for any $x,~{}x^{\prime}\in\mathbb{R}^{n}$,
@@
In the following, we provide the definitions of partitioned vector and partially strong convexity that will be used in this work.
@@
A global $\mu$-strongly convex function is also partially strongly convex since $\left(\nabla f(x)-\nabla f(x^{\prime})\right)^{\top}(x-x^{\prime})\geq\mu\|x-x^{\prime}\|^{2}\geq\mu\|x_{S}-x^{\prime}_{S}\|^{2}$.
Let $Q\in\mathbb{R}^{n\times n}$, $R\in\mathbb{R}^{n\times m}$ be an QR decomposition of $T^{\top}$, i.e., $Q$ is an orthogonal matrix and $R$ is an upper triangular matrix such that
@@
where $R_{1}\in\mathbb{R}^{m\times m}$, $Q_{1}\in\mathbb{R}^{n\times m}$ and $Q_{2}\in\mathbb{R}^{n\times(n-m)}$.
We relax the strong convexity requirement of $f(x)$ over the whole domain in the problem (1) to being strongly convex within the subspace satisfying the equality constraints. Let us define state $x^{\prime}$ by the linear transformation $x=Qx^{\prime}$ and consider the following assumption.
@@
When $m<n$, $g(x^{\prime})=f(Qx^{\prime})$ is $\mu$-partially strongly convex with respect to the partitioned vector $x^{\prime}_{S}$ with $S=\{n-m+1,n-m+2,\ldots,n\}$, namely, the vector $x^{\prime}$ is partitioned as $x^{\prime}=\left(x^{\prime}_{S^{c}},x^{\prime}_{S}\right)$ where $S^{c}=\{1,2,\ldots,n-m\}$.
@@ III Augmented Primal-dual algorithm
We define the augmented Lagrange for problem (1) by
@@
It is well-known that any equilibrium to the above dynamics satisfies the Karush–Kuhn–Tucker (KKT) conditions [21] and is thus an optimal solution to the problem under Assumptions 1, 2 and 3. In addition, exponential convergence of (4) is guaranteed when $f(x)$ is strongly convex. Nevertheless, we observe that when $f(x)=0$, the error system of (4) with respect to an equilibrium $(x^{*},\lambda^{*})$ becomes
@@
$A=\begin{bmatrix}-F&-T^{\top}\\ T&\mathbf{0}\end{bmatrix}$ (6)
where $F\in\mathbb{R}^{n\times n}$ and $T\in\mathbb{R}^{m\times n}$, $m\leq n$. If $F\succ 0$, and $T$ is of full row rank, then $A$ is Hurwitz.
@@
Since $A+A^{\top}\preceq 0$, all eigenvalues $s$ of $A$ have non-positive real parts. Let the non-zero vector $z=(x,\lambda)$ be an eigenvector of $A$ such that
@@
Assume that $\text{Re}(s)=0$, then $x=0$ since $F\succ 0$. Substituting it into (7), we have
$-T^{\top}\lambda=$ $0$ (8)
Recall that $T$ has full row rank, then (8) implies that $\lambda=0$, which contradicts that the eigenvector $z=(x,\lambda)\neq 0$. Therefore $\text{Re}(s)<0$, and thus $A$ is Hurwitz. ∎
@@
$\dot{x}=$ $-\nabla f(x)-T^{\top}\lambda$ (9a)
$\dot{\lambda}=$ $Tx-b.$ (9b)
@@ IV Exponential convergence by IQC
$\dot{z}=Az+Bu,\quad y=Cz,\quad u=\Delta(y)$ (11a)
$A=\begin{bmatrix}-T^{\top}T&-T^{\top}\\ T&\mathbf{0}\end{bmatrix},~{}B=\begin{bmatrix}-I\\ \mathbf{0}\end{bmatrix},~{}C=\begin{bmatrix}I&\mathbf{0}\end{bmatrix}$ (11b)
where $z=\left(\tilde{x},\tilde{\lambda}\right)$ and $u$ is nonlinearity $\Delta(y)$ added to the systems, given by
$u=\Delta(y):=\nabla f(y+x^{*})-\nabla f(x^{*}).$ (12)
In fact, $u=\nabla f(x)-\nabla(x^{*})$, and $y=x-x^{*}$. Since $f(x)$ is convex and $\nabla f(x)$ is $l$-Lipschitz, it satisfies the co-coercivity[24],
$\left(\nabla f(x)-\nabla f(x^{*})\right)^{\top}(x-x^{*})$
$\geq\frac{1}{l}\|\nabla f(x)-\nabla f(x^{*})\|^{2}.$ (13)
@@
$\begin{bmatrix}G_{\rho}(j\omega)\\ I\end{bmatrix}^{*}\Pi\begin{bmatrix}G_{\rho}(j\omega)\\ I\end{bmatrix}\preceq 0,~{}\text{for any~{}}\omega\in\mathbb{R}.$ (16)
@@
$-G_{\rho}(j\omega)-G_{\rho}^{\top}(-j\omega)\succeq-\frac{2}{l}I.$ (17)
which is equivalent to the system $(A_{\rho},B,-C)$ being input feedforward passive (IFP) with index $-\frac{2}{l}$ [26, 23].
Select the storage function $V_{s}=\frac{1}{2}\|x-x^{*}\|^{2}+\frac{1}{2}\|\lambda-\lambda^{*}\|^{2}\geq 0$, the derivative along the system $(A,B,-C)$ gives
$\dot{V}_{s}=-x^{\top}T^{\top}Tx+y^{\top}u$
meaning that the system $(A,B,-C)$ is passive, and $-G(j\omega)-G^{\top}(-j\omega)\succeq 0$ [26]. Since $A$ is Hurwitz by Lemma 1, $A_{\rho}$ is also Hurwitz for sufficiently small $\rho>0$. Thus, $G_{\rho}(j\omega)=G(j\omega-\rho)$ is analytical for all $\omega\in\mathbb{R}$ if $0\leq\rho<\textup{Re}(s_{\min}(-A))$, where $s_{\min}(\cdot)$ denotes the smallest eigenvalue. We conclude that there exists a sufficiently small $\rho>0$ such that (17) holds. Then, the system is $\rho$-exponential stable.
@@
Notice that $A^{\prime}$ is not Hurwitz. Nevertheless, recall from Assumption 4 that $g(x^{\prime})=f(Qx^{\prime})$ is partially strongly convex with respect to the partitioned vector $x^{\prime}_{S}$. Then, (18) can be rewritten as
@@
It can be similarly shown that (19) is passive using the same storage function $V_{s}$. Since an orthogonal linear transformation does not change the Lipschitz constant [4], by Assumptions 1, 2 and 4, $\Delta^{\prime}$ also satisfies the same IQC (14) with respect to $y^{\prime}$ and $u^{\prime}$. In addition, $\mathcal{A}$ is Hurwitz by Lemma 1. Thus, following similar arguments of the previous step, exponential convergence is guaranteed. ∎
@@
It is worth noting that the work [13] considers the problem
$\min_{x\in\mathbb{R}^{n},y\in\mathbb{R}^{m}}f(x)+g(y)\quad\text{subject to~{}}Ax+By=d$
@@ V Applications to Distributed Optimization
We apply the analysis of the equality-constrained problem in the previous section to distributed optimization to relax the requirement of strong convexity.
The distributed optimization problem $\min_{x\in\mathbb{R}^{n}}\sum_{i=1}^{N}f_{i}(x)$ can be rewritten into the standard form
$\min_{x_{i}\in\mathbb{R}^{n}}\sum_{i=1}^{N}f_{i}(x_{i}),~{}\text{subject to }\boldsymbol{L}x=\mathbf{0}$ (20)
@@
Under this assumption, (20) imposes that $x_{i}=x_{k}$, for all $i,k$, recovering the original distributed optimization problem. More on the distributed settings can be found in [27, 28], and references therein.
@@ V-A Relaxing local strong convexity
$\dot{x}=$ $-\nabla f(x)-\alpha\boldsymbol{L}x-\boldsymbol{L}\lambda$ (21a)
$\dot{\lambda}=$ $\boldsymbol{L}x$ (21b)
@@
$\min_{x^{\prime}_{i}\in\mathbb{R}^{n}}\sum_{i=1}^{N}f_{i}(\sum_{j=1}^{N}Q_{ij}x_{j}^{\prime}),$
$\text{subject to }\begin{bmatrix}\mathbf{\Lambda}&\mathbf{0}_{(N-1)n\times n}\end{bmatrix}x^{\prime}:=Tx^{\prime}=\mathbf{0}$ (22)
where $\mathbf{\Lambda}=\Lambda\otimes I_{n}$ and $T\in\mathbb{R}^{(N-1)n\times Nn}$ is of full row rank since $\Lambda$ is positive definite. Then, the analysis in the previous section can be applied to this problem. Recall that Assumption 4 requires $f(x)$ to be partially strongly convex only in the subspace that satisfies the equality constraint, i.e., when consensus is achieved among agents. Therefore, Assumption 4 is equivalent to saying that the sum $f(x)=\sum_{i=1}^{N}f_{i}(x)$ is strongly convex, which is a weaker condition than requiring strong convexity on local objective functions [16].

### v2

@@ Abstract
We show that the augmented primal-dual gradient algorithms can achieve global exponential convergence with partially strongly convex functions. In particular, the objective function only needs to be strongly convex in the subspace satisfying the equality constraint and can be generally convex elsewhere, provided the global Lipschitz condition for the gradient is satisfied. This condition implies that states outside the equality subspace will converge towards it exponentially fast. The analysis is then applied to distributed optimization, where the partially strong convexity can be relaxed to the restricted secant inequality condition, which is not necessarily convex. This work unifies global exponential convergence results for some existing centralized and distributed algorithms.
@@ I Introduction
The convergence of optimization algorithms has been an important research focus due to their wide range of applications. In particular, exponential convergence in continuous time, or linear convergence in discrete time, has attracted considerable attention recently. Researchers have proposed various conditions alternative to strong convexity of the objective function to ensure exponential/linear convergence, such as the Polyak-Łojasiewicz (PL) condition [1, 2], the error bound (EB) condition [3], the restricted secant inequality (RSI) condition [4], and the quadratic growth (QG) condition [5]. Notably, these conditions may not even require convexity [6].
Convergence analysis becomes more complicated when constraints are introduced. The augmented Lagrange multiplier method for handling equality constraints was introduced by [7, 8], and has been extensively studied since then [9]. Meanwhile, the exponential convergence has been of particular interest in recent years [10, 11, 12, 13, 14]. Specifically, [10] and [11] address this issue for strongly convex functions with linear equality and inequality constraints, respectively. For general nonlinear inequality constraints, only the so-called semi-global exponential stability is obtained without considering constraints to be linearly independent [12]. Currently, it remains uncertain whether a global exponential stability condition can be established for augmented primal-dual gradient dynamics in problems involving general nonlinear inequality constraints. The work in [14] extends exponential stability to non-smooth objective functions and general set constraints. Moreover, its application in distributed optimization is also extensively studied [15, 16, 17].
On the other hand, exponential convergence of primal-dual gradient algorithms by harnessing strong convexity from constraints is seldom considered. The most related work is [13], which addresses partially non-strongly convex problems but with decoupled objective functions and equality constraints. In this work, we aim to demonstrate explicitly how equality constraints can contribute to exponential convergence while relaxing the requirement for strong convexity. Our theoretical results are shown by utilizing the methods of integral quadratic constraints (IQCs) [18] and the Kalman-Yakubovich-Popov (KYP) lemma [19]. The proposed framework unifies and extends some existing results in both centralized and distributed optimization. Moreover, a comprehensive understanding of exponential convergence is provided in these contexts.
@@ II Preliminaries
$\min_{x\in\mathbb{R}^{n}}f(x)$ (1)
subject to $\text{subject to }Tx=b$
@@
$f(x)$ is $l$-Lipschitz smooth, i.e., there exists some constant $l>0$, such that for any $x,~{}x^{\prime}\in\mathbb{R}^{n}$,
@@
In the following, we provide the definitions of partitioned vector and partially strong convexity, which will be used in this work.
@@
A global $\mu$-strongly convex function is also partially strongly convex since
$\left(\nabla f(x)-\nabla f(x^{\prime})\right)^{\top}(x-x^{\prime})\geq\mu\|x-x^{\prime}\|^{2}\geq\mu\|x_{S}-x^{\prime}_{S}\|^{2}.$
Let $Q\in\mathbb{R}^{n\times n}$, $R\in\mathbb{R}^{n\times m}$ be a QR decomposition of $T^{\top}$, i.e., $Q$ is an orthogonal matrix and $R$ is an upper triangular matrix such that
@@
where $R_{1}\in\mathbb{R}^{m\times m}$, $Q_{1}\in\mathbb{R}^{n\times m}$ and $Q_{2}\in\mathbb{R}^{n\times(n-m)}$. We relax the strong convexity requirement of $f(x)$ over the whole domain in the problem (1) to being strongly convex within the subspace satisfying the equality constraints. Let us define state $x^{\prime}$ by the linear transformation $x=Qx^{\prime}$ and consider the following assumption.
@@
When $m<n$, $g(x^{\prime}):=f(Qx^{\prime})$ is $\mu$-partially strongly convex with respect to the partitioned vector $x^{\prime}_{S}$ with $S=\{n-m+1,n-m+2,\ldots,n\}$, namely, the vector $x^{\prime}$ is partitioned as $x^{\prime}=\left(x^{\prime}_{S^{c}},x^{\prime}_{S}\right)$ where $S^{c}=\{1,2,\ldots,n-m\}$.
@@ III Augmented Primal-dual algorithm
The augmented Lagrange for problem (1) is
@@
It is well-known that any equilibrium to the above dynamics satisfies the Karush–Kuhn–Tucker (KKT) conditions [21] and is thus an optimal solution to the problem under Assumptions 1, 2, and 3. In addition, exponential convergence of (4) is guaranteed when $f(x)$ is strongly convex. Nevertheless, we observe that when $f(x)=0$, the error system of (4) with respect to an equilibrium $(x^{*},\lambda^{*})$ becomes
@@
$A=\begin{bmatrix}-F&-T^{\top}\\ T&\mathbf{0}\end{bmatrix}$
where $F\in\mathbb{R}^{n\times n}$, $T\in\mathbb{R}^{m\times n}$, and $m\leq n$. If $F\succ 0$, and $T$ is of full row rank, then $A$ is Hurwitz.
@@
Since $A+A^{\top}\preceq 0$, all the eigenvalues $s$ of $A$ have non-positive real parts. Let the non-zero vector $z=(x,\lambda)$ be an eigenvector of $A$ such that
@@
Assume that $\text{Re}(s)=0$, then $x=0$ since $F\succ 0$. Substituting it into (6), we have $-T^{\top}\lambda=0$. Recall that $T$ has full row rank, then $T^{\top}\lambda=0$ implies that $\lambda=0$, which contradicts that the eigenvector $z=(x,\lambda)\neq 0$. Therefore, $\text{Re}(s)<0$, and thus $A$ is Hurwitz. ∎
@@
$\dot{x}=$ $-\nabla f(x)-T^{\top}\lambda$
$\dot{\lambda}=$ $Tx-b.$
@@ IV Exponential convergence by IQC
$\dot{z}=Az+Bu,\quad y=Cz,\quad u=\Delta(y)$ (8)
where $z=\left(\tilde{x},\tilde{\lambda}\right)$,
$A=\begin{bmatrix}-T^{\top}T&-T^{\top}\\ T&\mathbf{0}\end{bmatrix},B=\begin{bmatrix}-I\\ \mathbf{0}\end{bmatrix},C=\begin{bmatrix}I&\mathbf{0}\end{bmatrix}$
and $u$ is nonlinearity $\Delta(y)$ added to the systems, given by
$u=\Delta(y):=\nabla f(y+x^{*})-\nabla f(x^{*}),$
or equivalently, $u=\nabla f(x)-\nabla(x^{*})$, and $y=x-x^{*}$. Since $f(x)$ is convex and $\nabla f(x)$ is $l$-Lipschitz, it satisfies the co-coercivity[24], i.e.,
$\left(\nabla f(x)-\nabla f(x^{*})\right)^{\top}(x-x^{*})\geq\frac{1}{l}\|\nabla f(x)-\nabla f(x^{*})\|^{2}.$
@@
$\begin{bmatrix}G_{\rho}(-j\omega)\\ I\end{bmatrix}^{\top}\Pi\begin{bmatrix}G_{\rho}(j\omega)\\ I\end{bmatrix}\preceq 0,~{}\text{for all }\omega\in\mathbb{R}$ (11)
@@
$-G_{\rho}(j\omega)-G_{\rho}^{\top}(-j\omega)\succeq-\frac{2}{l}I,~{}\text{for all }\omega\in\mathbb{R}$ (12)
which is equivalent to system $(A_{\rho},B,-C)$ being input feedforward passive (IFP) with index $-\frac{2}{l}$ [26, 23]. Select the storage function
$V_{s}=\frac{1}{2}\|x-x^{*}\|^{2}+\frac{1}{2}\|\lambda-\lambda^{*}\|^{2}\geq 0,$
the derivative along the system $(A,B,-C)$ gives
$\dot{V}_{s}=-x^{\top}T^{\top}Tx+y^{\top}u,$
meaning that the system $(A,B,-C)$ is passive, and $-G(j\omega)-G^{\top}(-j\omega)\succeq 0$, for all $\omega\in\mathbb{R}$ [26].
Since $A$ is Hurwitz by Lemma 1, $G(s)$ has no poles for $\textup{Re}(s)>-\varepsilon$, i.e., it is analytic in the open vertical strip
$\left\{s\in\mathbb{C}\mid\textup{Re}(s)>-\varepsilon\right\}$
for some $\varepsilon>0$. Then $G(-s)$ is analytic for $\textup{Re}(-s)>-\varepsilon$. This implies that $G^{\top}(-s)$ is analytic for $\textup{Re}(s)<\varepsilon$. Therefore, the sum $G(s)+G^{\top}(-s)$ is analytic in the open vertical strip
$\left\{s\in\mathbb{C}\mid-\varepsilon<\textup{Re}(s)<\varepsilon\right\}.$
Recall that $G_{\rho}(j\omega)=G(j\omega-\rho)$, there exists a sufficiently small $\rho>0$ such that $G_{\rho}(j\omega)+G_{\rho}^{\top}(-j\omega)$ is analytic for all $\omega\in\mathbb{R}$. As it remains analytic as $\rho$ decreases to zero, and
$-\lim_{\rho\rightarrow 0}\left(G_{\rho}(j\omega)+G_{\rho}^{\top}(-j\omega)\right)=-G(j\omega)-G^{\top}(-j\omega)\succeq 0,$
for all $\omega\in\mathbb{R}$, we conclude that there exist sufficiently small $\delta\geq 0$ and $\rho(\delta)>0$ such that (12) holds, i.e.,
$-G_{\rho(\delta)}(j\omega)-G_{\rho(\delta)}^{\top}(-j\omega)\succeq-\delta I\succeq-\frac{2}{l}I,~{}\text{for all }\omega\in\mathbb{R}.$
Then, the system is $\rho(\delta)$-exponentially stable.
@@
Notice that $A^{\prime}$ is not Hurwitz. Nevertheless, recall from Assumption 4 that $g(x^{\prime})=f(Qx^{\prime})$ is $\mu$-partially strongly convex with respect to the partitioned vector $x^{\prime}_{S}$. Then, (13) can be rewritten as
@@
It can be similarly shown that (14) is passive. Since an orthogonal linear transformation does not change the Lipschitz constant [4], by Assumptions 1, 2 and 4, $\Delta^{\prime}$ also satisfies the same IQC (9) with respect to $y^{\prime}$ and $u^{\prime}$. In addition, $\mathcal{A}$ is Hurwitz by Lemma 1. Thus, following similar arguments to the previous step, exponential convergence is guaranteed. ∎
Remark 4
The proof above illustrates the convenience of incorporating a frequency-domain approach in the analysis, as it relies solely on a quadratic storage function rather than an explicit Lyapunov function to prove exponential stability. Specifically, we show that the Hurwitz property of $G(s)$ ensures that $G(j\omega)+G^{\top}(-j\omega)$ remains analytic under a small real shift. With this small perturbation, the system still preserves the essential IFP property needed for exponential stability. However, obtaining a tighter convergence rate may require the use of a more involved IQC than (9), which is out of the scope of this paper [24, 27, 28].
@@
It is worth noting that [13] considers the problem
$\min_{x\in\mathbb{R}^{n},y\in\mathbb{R}^{m}}f(x)+g(y)$
subject to $\text{subject to~{}}Ax+By=d$
@@ V Applications to Distributed Optimization
We apply the analysis of the equality-constrained problem in the previous section to distributed optimization to relax the requirement of strong convexity. The distributed optimization problem is
$\min_{x\in\mathbb{R}^{n}}\sum_{i=1}^{N}f_{i}(x)$
which can be rewritten as the standard form
$\min_{x_{i}\in\mathbb{R}^{n},~{}i=1,\ldots,N}\sum_{i=1}^{N}f_{i}(x_{i})$ (15)
subject to $\qquad\text{subject to }\boldsymbol{L}x=\mathbf{0}$
@@
Under this assumption, (15) imposes that $x_{i}=x_{k}$, for all $i,k$, recovering the original distributed optimization problem. More on the distributed settings can be found in [29, 30] and references therein.
@@ V-A Relaxing local strong convexity
$\dot{x}=$ $-\nabla f(x)-\alpha\boldsymbol{L}x-\boldsymbol{L}\lambda$ (16)
$\dot{\lambda}=$ $\boldsymbol{L}x$
@@
$\min_{x^{\prime}_{i}\in\mathbb{R}^{n},~{}i=1,\ldots,N}\sum_{i=1}^{N}f_{i}(\sum_{j=1}^{N}Q_{ij}x_{j}^{\prime}),$
$\text{subject to }\begin{bmatrix}\mathbf{\Lambda}&\mathbf{0}_{(N-1)n\times n}\end{bmatrix}x^{\prime}=:\mathbf{T}x^{\prime}=\mathbf{0}$
where $\mathbf{\Lambda}=\Lambda\otimes I_{n}$ and $\mathbf{T}\in\mathbb{R}^{(N-1)n\times Nn}$ is of full row rank since $\Lambda$ is positive definite. Then, the analysis in the previous section can be applied to this problem. Recall that Assumption 4 requires $f(x)$ to be partially strongly convex only in the subspace that satisfies the equality constraint, i.e., when consensus is achieved among agents. Therefore, Assumption 4 is equivalent to saying that the sum $f(x)=\sum_{i=1}^{N}f_{i}(x)$ is strongly convex, which is a weaker condition than requiring strong convexity on local objective functions [16].